The total market value of Quinlan Carpet's assets is $1.2 Million. The total market value of all of their debt is $450,000. The capital structure of the company is made up of only debt and common equity. There are 50,000 common shares outstanding. The company has declared a 20% stock dividend, with the ex-dividend date being tomorrow.

Answer the following:

All else equal, what do you expect the ex-dividend price to be? Price = $ (Round your answer to 2 decimal places. Do NOT use commas (eg. 43.21)

All else equal, what will be the total ex-dividend market value of the company's equity?

Market Value of Equity (ex-dividend) = $ (Round your answer to the closest dollar. Do NOT use commas (i.e. 1234567)
Answer #1

current price = (assets-debt)/shares outstanding

=(1200000-450000)/50000=15

price after stock dividend = current price/stock dividend ratio

=15/(1+0.2)

=12.5

share after stock dividend = current shares*stock dividend ratio = 50000*1.2=60000

market value = price after*shares after = 12.5*60000=750000